OpenStudy (anonymous):

will medal for best awnser. How many triangles can be constructed with side lengths of 7 cm, 10 cm, and 18 cm? A. 0 B. 1 C. 2 D. an infinite number

OpenStudy (anonymous):

@mathsciencehistory

OpenStudy (anonymous):

@bibby @inowalst @undeadknight26 @welshfella @Data_LG2

OpenStudy (anonymous):

what do you think it is

OpenStudy (anonymous):

i believe B

OpenStudy (welshfella):

for a triangle to be drawn the sum of any 2 sides must be greater than the third side do the values in your question numbers fit this requirement/

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so a?

OpenStudy (anonymous):

im not sure

OpenStudy (welshfella):

17 + 10 = 17 the other side is 18

OpenStudy (anonymous):

welshfella , do you mean that a triangle cannot have three sides with 2cm each

OpenStudy (welshfella):

No 2+2 > 2 so that is a triangle

OpenStudy (anonymous):

wait what?

OpenStudy (welshfella):

what?

OpenStudy (anonymous):

i dont understnad

OpenStudy (welshfella):

oops i should have said 7 + 10 not 17 + 10!! 2 sides are 7 and 10 7 + 10 = 17 - and the 3rd side is 18

OpenStudy (anonymous):

sorry someone beat you to it but thanks for the help its 0

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

that's ok - you understand the concept, though?